OpenStudy (anonymous):

(a*11.34+9.15) /(a+1)=10.61 need steps I know the answer

sam (.sam.):

(11.34a+9.15)/(a+1)=10.61 (11.34a+9.15)/(10.61)=a+1 0.0943(11.34a+9.15)=a+1 1.0688a+0.8624=a+1 0.0688a+0.8624=1 0.0688a=0.1376 a=2

OpenStudy (anonymous):

Thanks Sam got to figure out the steps you done but I think I can do that

OpenStudy (anonymous):

How did you get rid of the a on the right on the 4th step

OpenStudy (anonymous):

if it was me, i would clear the decimals first and start with \(\frac{a11.34+9.15}{a+1}=10.61\) \[\frac{1134a+915}{a+1}=1061\] all the work will be the same, but the annoying decimals will be gone

OpenStudy (anonymous):

\[\frac{1134a+915}{a+1}=1061\]\[1134a+915=1061(a+1)\] \[1134a+915=1061a+1061\] etc

OpenStudy (anonymous):

subtract \(1061a\) from both sides, subtract 915 from both sides and then divide

sam (.sam.):

Here 1.0688a+0.8624=a+1 to here? 0.0688a+0.8624=1 -------------------------------------- substract both sides by "a", 1.0688a-a+0.8624=a-a+1 0.0688a+0.8624=1

OpenStudy (anonymous):

Thanks Sam. satellite73 your answer worked if I do not remove the decimals